# Stability analysis through folds: An end-loaded elastica with a lever arm

### v3

@@ 4.1 Varying Parameter in Free end
In the first section of the examples, the parameters associated with the free end $(s=l)$ are varied. These include the magnitude of the tip load $P$, direction of the tip load $\alpha$, length of the load arm $\epsilon$, and the orientation of the arm $\psi$. Moreover, the parameters $P$ and $\epsilon$ can assume negative values, as they represent the opposite direction of the force vector $\mathbf{F}$ and the arm vector $\Psi$ respectively. In the presented examples, stiffness $K$ and length $l$ are set to unit values and, therefore, $P=|\mathbf{F}|$ and $\epsilon=\Delta$.
@@ 4.1.1 Rotating Arm
We begin with a numerical example in which continuation is performed with respect to $\psi$, simulating the quasi-static rotation of the arm. The system is $2\pi$-periodic in $\psi$, i.e., the system corresponding to $\psi=k$ and $\psi=2\pi+k$ exhibit identical features for any real $k$. Consequently, the solution family also exhibits $2\pi$- periodic behavior. The analysis begins with parameters $P=\pi^{2}/4$ and $\epsilon=0.25$, and continuation is initiated from the straight trivial equilibrium at $\psi=0$ which has an index of one. When $\psi$ is the varying parameter, the ordinate in the distinguished bifurcation diagrams, as specified by (29) is
@@
$=\left(\Psi(l)\times\mathbf{F}\right)\cdot\mathbf{e}_{z}=\mathbf{m}(l)\cdot\mathbf{e}_{z}.$
@@
Figure 4: The distinguished bifurcation diagram for arm $\epsilon=0.25$ rotating for loads (a)$P=\frac{\pi^{2}}{4}$ and (b) $P=\frac{9\pi^{2}}{4}$. The intermediate configurations along the family of stable equilibrium are labeled and depicted at the top. The stable configurations before and after the snap-back instability are also displayed. In (b), all equilibria are unstable. Although the folds are present no information on snapping behavior can be inferred, as no stable equilibrium exists. The tip trace in the foldless region is shown (in green).
In the last step, the boundary condition (38b) is applied. This ordinate $(\mathbf{m}(l).\mathbf{e}_{z})$ is evaluated from the numerical continuation solutions and plotted against the parameter $\psi$ to generate the associated distinguished bifurcation diagram, as shown in Figure 4. This plot exhibits folds, indicating stability transitions. The direction of change in index near folds is inferred from Figure 1(b) and is indicated in Figures 4. The dynamical aspects of the snapping motion are not discussed in this paper, and for further details, refer [40, 18]. The solutions and their projections exhibit $2\pi$- periodic characteristics. Figure 4(a) also displays the elastica configurations at intermediate values of $\psi$. The potential snapping motion depicting the initial and final configurations is also shown. This motion depends on the current state and the direction of current parameter change, and the snap jump is indicated by a red arrow.
Figure 4(b) displays the bifurcation diagrams for a family of equilibria at a higher load, $P=9\pi^{2}/4$. Under this loading, the straight trivial equilibrium has an index of two (Section 3.1). These plots also exhibit folds, which, although appear qualitatively similar to the case of $P=\pi^{2}/4$ but differ significantly in their stability characteristics. In this scenario, only families of equilibria with indices one and two exist, corresponding to unstable equilibria. To obtain a family of stable equilibria, the continuation process must be initiated from a different solution. The potential snap-back instability can be assessed only when a family of stable equilibria exists.
@@
Figure 5: The distinguished bifurcation diagram for an arm $\epsilon=0.5$ rotating for loads (a) $P=\frac{\pi^{2}}{4}$ and (b)$P=\frac{9\pi^{2}}{4}$. Intermediate configurations along this family are labeled and depicted at the top. Additionally, the stable configurations before and after the snap-back instability are displayed. The trajectory of the arm’s tip preceding the snap-back instability is indicated in green
A similar analysis is performed for a longer arm $\epsilon=0.5$, and Figure 5 displays the corresponding plots. The straight trivial solutions with this arm have the same index as those of $\epsilon=0.25$, i.e., one and two for loads $P=\pi^{2}/4$ and $P=9\pi^{2}/4$, respectively. The bifurcation diagram and the configurations appear qualitatively similar to those in the case of $\epsilon=0.25$. However, the higher load case of $P=9\pi^{2}/4$ exhibits significant differences. Here, we observe several folds, with more than two equilibria existing for an identical value of $\psi$ suggesting the presence of multistability. Unlike the previous example for $P=9\pi^{2}/4$ and $\epsilon=0.25$ in Figure 4(b), a family of equilibria with index zero exists here. Consequently, snap-back instability occurs as $\psi$ is varied beyond the folds. The plots also illustrate several equilibrium configurations during this maneuver. Furthermore, the stable equilibria before and after the set of snap-back instability are also displayed. We assume that the geometry of the lever arm and elastica allows for self-intersection, and the resulting configurations as displayed in the plots are valid. For a zero arm, when the load $P$ exceeds the critical buckling length, it buckles into a non-straight configuration, deflecting either to the right or to the left, with both configurations being mirror images of each other. The non-zero asymmetrical arm also results in similar post-buckling structures, but they are not mirror images. The arm acts as an imperfect parameter breaking the symmetry. Furthermore, the arm also induces bi-stability in the buckled configurations, meaning there is more than one stable equilibria for some values of $\psi$ in buckled states, as shown in Figure 5(b). Subsequently, more than two stable equilibria exist in some instances. When $\psi$ is varied past the fold, snap-back instability arises and leads to one of the available stable equilibria. The specific equilibrium reached depends on the amount of energy released during snapping and the dissipation of the system and is beyond the scope of this paper. With appropriate technical considerations, novel snapping, triggering, and switching mechanisms can be realized. The periodicity of the system with respect to $\psi$ allows the reproducibility of the snap-back instability, i.e., the similar snapping motion can be replicated by tuning the parameter forward by $2\pi$. Furthermore, snapping motion can be generated by reversing the parameter $\psi$, but occurs at a fold in the reverse direction.
@@
Figure 6: (a) The bifurcation diagram from the previous analysis adjusted between 00 and $2\pi$. Equilibria are labeled according to the order which continuation crosses a specified $\psi$. (b) Distinguished bifurcation diagrams for $\epsilon$ varying from $0.5$ to $-0.5$ at $\psi=0$, showing the presence of folds and bifurcations. The equilibria corresponding to the given labels are shown at the top. The potential snap-back instability is also illustrated.
4.1.2 Varying Arm Length
We continue this analysis by varying the parameter associated with the arm’s length $\epsilon$. The equilibria with $\epsilon=0.5$ at $\psi=0$, $\psi=\pi/12$, $\psi=\pi/6$, and $\psi=\pi/2$ from the previous analysis are chosen as initial solutions, and continuation is performed along $\epsilon$ from $0.5$ to $-0.5$. Figure 6(a) displays the bifurcation plot from the previous analysis, truncated between 00 and $2\pi$ to obtain all possible equilibria at a given $\psi$. Owing to the $2\pi$-periodicity of the system, the section of the plot between $4\pi$ and $6\pi$ is represented within the interval 00 to $2\pi$. In this case, the ordinate of the distinguished bifurcation diagram is
which is the force component along the arm vector. Figure 6(b) and Figure 7 display the $\beta$ vs. $\epsilon$ plots associated with various $\psi$. A few plots exhibit folds and stability information is encoded according to Figure 1(b). Let us first consider the symmetric case of $\psi=0$. Here, the ordinate $\beta$, which is the component of force along the arm vector, has identical values for deflections to both the right and left, resulting in overlapping plots. Consequently, the supercritical pitchfork bifurcation diagram appears only as a half plot. The plots separate as $\psi$ is varied to a non-zero value. The half-pitchfork plot for $\psi=0$, separates into two curves for $\psi=\frac{\pi}{12}$, resembling an imperfect system. The parameter $\psi$ acts as an imperfection parameter perturbing the the perfect system obtained for $\psi=0$. Interestingly, the equilibria chosen as starting solutions lie on branches of solutions continued from other initial solutions. In these instances, the change in the indices aligns with the observed folds. If a stable equilibrium exists, it transitions to the other stable equilibrium lying on the other curve, when the parameter is varied past the folds. Based on this observation, we illustrate a few potential configurations that may result from the snapping motion. In each case, multiple snapping configurations are possible, and the exact configuration to which the system reaches depends on the energy released during snapping and the associated dissipation.
@@ 4.1.2 Varying Arm Length
Figure 7: Distinguished bifurcation diagrams for the arm length $\epsilon$ varying from from $0.5$ to $-0.5$ at arm angles (a) $\psi=\pi/12$, (b) $\psi=\pi/6$, and (c) $\psi=\pi/2$. Several instances of snap-back instability can be noticed and a few of them are illustrated. An enlarged view of a section of plot (c) is provided to clearly illustrate the folds.
@@ 4.1.3 Rotating Load
$\begin{split}\frac{\partial B}{\partial\alpha}&=-\frac{\partial}{\partial\alpha}\left(\left(P\sin\alpha\mathbf{e}_{x}-P\cos\alpha\mathbf{e}_{y}\right)\cdot\left(\mathbf{r}(l)+\Psi(l)\right)\right),\\ &=\mathbf{e}_{z}\cdot\left(\left(\mathbf{r}(l)+\Psi(l)\right)\times\mathbf{n}(l)\right),\\ &=\mathbf{e}_{z}\cdot\left(\mathbf{m}(l)+\mathbf{r}(l)\times\mathbf{n}(l)\right).\end{split}$ (48)
(a) $P=\frac{\pi^{2}}{4}$
(b) $P=\frac{9\pi^{2}}{4}$
Figure 8: Distinguished Bifurcation Diagrams when load $P$ with an arm $\epsilon=0.5$ changes its orientation, making full rotations for $P=\frac{\pi^{2}}{4}$ and $P=\frac{9\pi^{2}}{4}$. The intermediate configurations along this stable-family are displayed at the top. The tip trace before the snap-back instability is also displayed (in green).
@@
Figure 9: (a) The bifurcation diagram from the previous analysis, adjusted between 00 and $2\pi$. The equilibria are labeled based on the order in which the continuation crosses a specified value of $\alpha$. Continuation is then performed from these labeled equilibria along $P$. (b) The bifurcation diagram when continuation is performed from the equilibria corresponding to $\alpha=0$. Folds appear in the diagram, indicating snap-back instability. The labeled equilibria and the snap-back instabilities between the equilibria are displayed both at the top and on the side of the plot.
@@
Figure 10: The bifurcation diagram when continuation is performed from the equilibria corresponding to $\alpha=0$. Folds appear in the diagram, indicating the presence of snap-back instability. The labeled equilibria and the snap-back instability occurring between the equilibria are shown at the top and on the side.
@@ 4.1.4 Varying Load Magnitude
We now examine the effect of load’s magnitude $P$ on the stability of our elastica setup. We employ the configurations from the previous example as starting solutions and perform continuation along $P$ from $49\pi^{2}/$ 4 to $-9\pi^{2}/4$. The ordinate of the corresponding distinguished bifurcation diagrams is
@@
which is the height of the point of application of force measured along the force vector $\mathbf{F}$. Figure 10 displays the response of the family of systems for this maneuver through bifurcation plots. The equilibria chosen as starting solutions lie on the continuation branches that originate from other initial solutions. The changes in the index are consistent with the predictions (Figure 1(b)). These diagrams also resemble the unfolding of perfect systems. These equilibrium paths collapse onto the equilibrium path of the perfect system, shown in Figure 6b, as the parameters $\psi$ and $\alpha$ approach zero. We can also notice the response of the branches emanating from the solutions with index $2$. The response is not reversible and reproducible with solely respect to $P$. However, a combination of parameters $\psi$, $\alpha$ and $P$ can be employed to replicate the behavior. This scenario can be employed in switches, where loads due to gravity, electric field or magnetic field, are varied.
@@
Figure 11: The bifurcation plots as the parameter $\theta_{o}$ is varied. Some instances of stable equilibria are labeled and displayed along with the potential snap-back instability. The snap-back instability cases for $P=\frac{9\pi^{2}}{4}$ are shown in Figure 12. The tip trace before the snap-back instability is shown in green.
4.2 Varying Parameter in Fixed end
Finally, we focus on the stability transitions when the varying parameter appears at the fixed end $s=0$. In this case, we have only one parameter that is the rotation of the clamped end, $\theta_{o}$.
4.2.1 Rotating base
The analysis is performed for a non-zero arm $\Delta=0.5$ held at $\psi=\pi/2$ for two loads $P=\pi^{2}/4$ and $49\pi^{2}/$ 4. In this case, the ordinate of the bifurcation diagrams given by (25) is
$\left[\frac{\partial\mathcal{L}}{\partial\zeta^{\prime}}\cdot\frac{\partial\zeta}{\partial\theta_{o}}\right]_{s=0}=K\theta^{\prime}(0)\equiv\mathbf{m}(0)\cdot\mathbf{e}_{z},$ (50)
@@ 4.2.1 Rotating base
$\mathbf{m}+\mathbf{r}\times\mathbf{n}$ (51)
@@
$\mathbf{m}(0)=\mathbf{m}(l)+\mathbf{r}(l)\times\mathbf{n}(l).$ (52)
In our setup, $\alpha$ and $\theta_{o}$ are related through $\alpha=-\theta_{o}$. As a result, the ordinate in the current case is same as that presented in subsection 4.1.3 of section 4.1, while the abscissa is sign-reversed. The plots in Figure 11 represent horizontally mirrored versions of those shown in Figure 8. The evolution of elastica configurations as the clamped end is rotated, are also indicated in the plots, along with the snap-back instability, illustrating the catapult-like behavior. The self-contact within the elastica and with the lever arm is disregarded. The lower load case of $\pi^{2}/4$ led to equilibria that don’t involve self-contact. The higher load case of $49\pi^{2}/$ 4 generated multi-stable equilibria that formed loops and the snap-back instability would result in more than one equilibrium. Two families of stable equilibrium can be observed. The equilibrium configurations of one family are illustrated in Figure 11. The configurations for the second family, along with the snap-back instability, are shown in Figure 12.

### v5

@@ 4.1 Varying Parameter in Free end
In the first subsection, the parameters associated with the free end $(s=l)$ are varied. These include the magnitude of the tip load $P$, direction of the tip load $\alpha$, length of the load arm $\epsilon$, and the orientation of the arm $\psi$. Moreover, the parameters $P$ and $\epsilon$ can assume negative values, and they indicate the negative direction of the force vector $\mathbf{F}$ and the arm vector $\boldsymbol{\Psi}$ respectively. In the presented examples, stiffness $K$ and length $l$ are set to unit values and, therefore $P=|\mathbf{F}|$ and $\epsilon=\Delta$.
@@ 4.1.1 Rotating Arm
We begin with a numerical example in which continuation is performed with respect to $\psi$, simulating the quasi-static rotation of the arm. The system is $2\pi$-periodic in $\psi$, i.e., the system corresponding to $\psi=k$ and $\psi=2\pi+k$ exhibit identical features for any real $k$. Consequently, the solution family also exhibits $2\pi$- periodic behavior. The analysis begins with parameters $P=\pi^{2}/4$ and $\epsilon=0.25$, and continuation is initiated from the straight trivial equilibrium at $\psi=0$ which has an index of one ($P_{cr,1}^{+}<P<P_{cr,2}^{+}$ in Section 3.2). When $\psi$ is the varying parameter, the ordinate in the distinguished bifurcation diagrams, as specified by (29) is
@@
$=\left(\boldsymbol{\Psi}(l)\times\mathbf{F}\right)\cdot\mathbf{e}_{z}=\mathbf{m}(l)\cdot\mathbf{e}_{z}.$
@@
[…]
In the last step, the boundary condition (38b) is applied. This ordinate $\mathbf{m}(l).\mathbf{e}_{z}$ is evaluated from the numerical continuation solutions and plotted against the parameter $\psi$ to generate the associated distinguished bifurcation diagram, as shown in Figure 4. This plot exhibits folds, indicating stability transitions. The direction of change in index near folds is inferred from Figure 1(b) and is indicated in Figures 4. The equilibrium at $\psi=\pi$ is equivalent to the case of $\epsilon<0$ in Section 3.2, where $P<P_{cr,1}^{-}$ corresponds to index zero and agrees with the diagrams. The dynamical aspects of the snapping motion are not discussed in this paper, and for further details, refer to [40, 18]. The solutions and their projections exhibit $2\pi$- periodic characteristics. Figure 4(a) also displays the elastica configurations at intermediate values of $\psi$. The potential snapping motion depicting the initial and final configurations is also shown. This motion depends on the current state and the direction of current parameter change, and the snap jump is indicated by a red arrow.
Figure 4(b) displays the bifurcation diagrams for a family of equilibria at a higher load, $P=9\pi^{2}/4$. Under this loading, the straight trivial equilibrium has an index of two ($P_{cr,2}^{+}<P<P_{cr,3}^{+}$ in Section 3.2). These plots also exhibit folds, which, although appear qualitatively similar to the case of $P=\pi^{2}/4$ but differ significantly in their stability characteristics. In this scenario, only families of equilibria with indices one and two exist, corresponding to unstable equilibria. To obtain a family of stable equilibria, the continuation process must be initiated from a different solution. The potential snap-back instability can be assessed only when a family of stable equilibria exists.
@@
[…]
@@
Figure 6: (a) The bifurcation diagram from the previous analysis adjusted between 0 and $2\pi$. Equilibria are labeled according to the order which continuation traverses the specified $\psi$. (b) Distinguished bifurcation diagrams for $\epsilon$ varying from $0.5$ to 0.5-0.5 at $\psi=0$, showing the presence of folds and bifurcations. The equilibria corresponding to the given labels are shown at the top. The potential snap-back instability is also illustrated.
A similar analysis is performed for a longer arm $\epsilon=0.5$, and Figure 5 displays the corresponding plots. The straight trivial solutions with this arm have the same index as those of $\epsilon=0.25$ i.e, one and two for loads $P=\pi^{2}/4$ and $P=9\pi^{2}/4$, respectively. The bifurcation diagram and the configurations for $P=\pi^{2}/4$, shown in Figure 5(a) appear qualitatively similar to those in the case of $\epsilon=0.25$ displayed in Figure 4(a). However, the higher load case of $P=9\pi^{2}/4$ exhibits significant differences. Here, we observe several folds, with more than two equilibria existing for an identical value of $\psi$ suggesting the presence of multi-stability. Unlike the previous example for $P=9\pi^{2}/4$ and $\epsilon=0.25$ in Figure 4(b), a family of equilibria with index zero exists here. Consequently, snap-back instability occurs as $\psi$ is varied beyond the folds. The plots also illustrate several equilibrium configurations during this maneuver. Furthermore, the stable equilibria before and after the set of snap-back instability are also displayed. We assume that the geometry of the lever arm and elastica allows for self-intersection, and the resulting configurations as displayed in the plots are valid. For a zero arm, when the load $P$ exceeds the critical buckling length, it buckles into a non-straight configuration, deflecting either to the right or to the left, with both configurations being mirror images of each other. The non-zero asymmetrical arm also results in similar post-buckling structures, but they are not mirror images. The arm acts as an imperfect parameter breaking the symmetry. Furthermore, the arm also induces bi-stability in the buckled configurations, meaning there is more than one stable equilibria for some values of $\psi$ in buckled states, as shown in Figure 5(b). Subsequently, more than two stable equilibria exist in some instances. When $\psi$ is varied past the fold, snap-back instability arises and leads to one of the available stable equilibria. The specific equilibrium reached depends on the amount of energy released during snapping and the dissipation of the system and is beyond the scope of this paper. With appropriate technical considerations, novel snapping, triggering, and switching mechanisms can be realized. The periodicity of the system with respect to $\psi$ allows the reproducibility of the snap-back instability, i.e., the similar snapping motion can be replicated by tuning the parameter forward by $2\pi$. Furthermore, snapping motion can be generated by reversing the parameter $\psi$, but occurs at a fold in the reverse direction.
@@
Figure 7: Distinguished bifurcation diagrams for the arm length $\epsilon$ varying from from $0.5$ to 0.5-0.5 at arm angles (a) $\psi=\pi/12$, (b) $\psi=\pi/6$, and (c) $\psi=\pi/2$. Several instances of snap-back instability can be noticed and a few of them are illustrated. An enlarged view of a section of plot (c) is provided to clearly illustrate the folds.
(a) $P=\frac{\pi^{2}}{4}$
(b) $P=\frac{9\pi^{2}}{4}$
Figure 8: Distinguished Bifurcation Diagrams when load $P$ with an arm $\epsilon=0.5$ varies its direction $\alpha$, completing several full rotations for $P=\frac{\pi^{2}}{4}$ and $P=\frac{9\pi^{2}}{4}$. The intermediate configurations along this stable-family are displayed at the top. The tip trace before the snap-back instability is also displayed (in green).
4.1.2 Varying Arm Length
We continue this analysis by varying the parameter associated with the arm’s length $\epsilon$. The equilibria with $\epsilon=0.5$ at $\psi=0$, $\psi=\pi/12$, $\psi=\pi/6$, and $\psi=\pi/2$ from the previous analysis are chosen as initial solutions, and continuation is performed along $\epsilon$ from $0.5$ to 0.5-0.5. Figure 6(a) displays the bifurcation plot from the previous analysis, truncated between 0 and $2\pi$ to obtain all possible equilibria at a given $\psi$. Owing to the $2\pi$-periodicity of the system, the section of the plot between $4\pi$ and $6\pi$ is represented within the interval 0 to $2\pi$. In this case, the ordinate of the distinguished bifurcation diagram is
which is the force component along the arm vector. Figure 6(b) and Figure 7 display the $\beta$ vs. $\epsilon$ plots associated with various $\psi$. A few plots exhibit folds and stability information is encoded according to Figure 1(b). Let us first consider the symmetric case of $\psi=0$. Here, the ordinate $\beta$, the load vector’s projection along the arm $\Psi$, has identical values for deflections to both the right and left, resulting in overlapping plots. Consequently, the supercritical pitchfork bifurcation diagram appears only as a half plot, which splits as $\psi$ is varied to a non-zero value. The half-pitchfork plot for $\psi=0$, separates into two curves for $\psi=\frac{\pi}{12}$, resembling an imperfect system. The parameter $\psi$ acts as an imperfection parameter perturbing the the perfect system obtained for $\psi=0$. Interestingly, the equilibria chosen as starting solutions lie on branches of solutions continued from other initial solutions. In these instances, the change in the indices is consistent with the observed folds. If a stable equilibrium exists, it transitions to the other stable equilibrium lying on the other curve, when the parameter is varied past the folds. Based on this observation, we illustrate a few potential configurations that may result from the snapping motion. In each case, multiple snapping configurations are possible, and the exact configuration to which the system reaches depends on the energy released during snapping and the associated dissipation.
@@ 4.1.3 Rotating Load
Figure 9: (a) The bifurcation diagram from the previous analysis, adjusted between 0 and $2\pi$. The equilibria are labeled based on the order in which the continuation crosses a specified value of $\alpha$. Continuation is then performed from these labeled equilibria along $P$. (b) The bifurcation diagram when continuation is performed from the equilibria corresponding to $\alpha=0$. The labeled equilibria are displayed both at the top of the plot.
@@
Figure 10: The bifurcation diagram when continuation is performed from the equilibria corresponding to $\alpha=0$. The labeled equilibria and the snap-back instability occurring between the equilibria are shown at the top and on the side.
$\begin{split}\frac{\partial B}{\partial\alpha}&=-\frac{\partial}{\partial\alpha}\left(\left(P\sin\alpha\mathbf{e}_{x}-P\cos\alpha\mathbf{e}_{y}\right)\cdot\left(\mathbf{r}(l)+\boldsymbol{\Psi}(l)\right)\right),\\ &=\left(\left(\mathbf{r}(l)+\boldsymbol{\Psi}(l)\right)\times\mathbf{n}(l)\right)\cdot\mathbf{e}_{z},\\ &=\left(\mathbf{m}(l)+\mathbf{r}(l)\times\mathbf{n}(l)\right)\cdot\mathbf{e}_{z}.\end{split}$ (48)
@@ 4.1.4 Varying Load Magnitude
We now examine the effect of load’s magnitude $P$ on the stability of our elastica setup. We employ the configurations from the previous example as starting solutions and perform continuation along $P$ from $9\pi^{2}/4$ to $9\pi^{2}/4$. The ordinate of the corresponding distinguished bifurcation diagrams is
@@
which is the height of the point of application of force projected on the load vector $\mathbf{F}$. Figure 9(a) displays the bifurcation plot from the previous analysis (subsection 4.1.3 Figure 8(b)), truncated between 0 and $2\pi$. Since the system is $2\pi$-periodic with respect to $\alpha$, the section of the plot between $2\pi$ and $4\pi$ is displayed to include all possible equilibria. Note that the section between 0 and $2\pi$ includes only the equilibria associated with forward continuation. Since backward continuation is not performed in this analysis, the corresponding equilibria are absent. Figure 9(b) and Figure 10 displays the response of the elastica system for this maneuver through bifurcation plots. The equilibria chosen as starting solutions lie on the continuation branches that originate from other initial solutions. The changes in the index are consistent with the predictions (Figure 1(b)). These diagrams also resemble the unfolding of perfect systems. These equilibrium paths collapse onto the equilibrium path of the perfect system, shown in Figure 6b, as the parameters $\psi$ and $\alpha$ approach zero. We can also notice the response of the branches emanating from the solutions with index $2$. The response is not reversible and reproducible with solely respect to $P$. However, a combination of parameters $\psi$, $\alpha$ and $P$ can be employed to replicate the behavior. This scenario can be employed in switches, where load fields due to gravity, electric field or magnetic field, can be tuned.
4.2 Varying Parameter in Fixed end
Finally, we focus on the stability transitions when the varying parameter appears at the fixed end $s=0$. In this case, we have only one parameter that is the rotation of the clamped end $\theta_{o}$.
4.2.1 Rotating base
The analysis is performed for a non-zero arm $\Delta=0.5$ held at $\psi=\pi/2$ for two loads $P=\pi^{2}/4$ and $9\pi^{2}/4$. In this case, the ordinate of the bifurcation diagrams given by (25) is
$\left[\frac{\partial\mathcal{L}}{\partial\zeta^{\prime}}\cdot\frac{\partial\zeta}{\partial\theta_{o}}\right]_{s=0}$ $=K\theta^{\prime}(0)$
$\equiv\mathbf{m}(0)\cdot\mathbf{e}_{z},$
@@ 4.2.1 Rotating base
Figure 11: The bifurcation plots as the clamped end $\theta_{o}$ is varied. Selected cases of stable equilibria are labeled and displayed along with the potential snap-back instability. The snap-back instability cases for $P=\frac{9\pi^{2}}{4}$ are shown in Figure 12. The tip trace before the snap-back instability is shown in green.
@@
$\mathbf{m}+\mathbf{r}\times\mathbf{n}$
@@
$\mathbf{m}(0)=\mathbf{m}(l)+\mathbf{r}(l)\times\mathbf{n}(l).$
In our setup, the parameters $\alpha$ and $\theta_{o}$ are interdependent; varying $\alpha$ is equivalent to varying $\theta_{o}$ in the negative direction. As a result, the ordinate in the current case is same as that presented in subsection 4.1.3 of section 4.1, while the abscissa is sign-reversed. The plots in Figure 11 represent horizontally mirrored versions of those shown in Figure 8. The evolution of elastica configurations as the clamped end is rotated, are also indicated in the plots, along with the snap-back instability, illustrating the catapult-like behavior. The self-contact within the elastica and with the lever arm is disregarded. The lower load case of $\pi^{2}/4$ led to equilibria that don’t involve self-contact. The higher load case of $9\pi^{2}/4$ generated multi-stable equilibria that formed loops and the snap-back instability would result in more than one equilibrium. Two families of stable equilibrium can be observed. The equilibrium configurations of one family are illustrated in Figure 11. The configurations for the second family, along with the snap-back instability, are shown in Figure 12.